OpenStudy (anonymous):

What is the equation of the line that passes through (4, -1) and (-2, 3)? a) 2x-3y-5=0 b) 2x+3y-5=0 c) -2x+3y-5=0

pooja195 (pooja195):

We need to find the slope first so we need to use the slope formula \[\huge~\rm~\frac{ y_2-y_1 }{ x_2-x_1 }\]

OpenStudy (pawanyadav):

y-y1={y2-y1} /(x2-x1) ×(x-x1)

OpenStudy (anonymous):

I got 1/3 for the slope formula, is that right?? @pooja195

pooja195 (pooja195):

i was thinking 1 .. ...

pooja195 (pooja195):

3-1=2 -2+4=2 2/2=1

OpenStudy (pawanyadav):

Put y2=3 Y1=-1 x2=-2 x1=4

pooja195 (pooja195):

???

OpenStudy (anonymous):

i did the formula wrong... i got it now

pooja195 (pooja195):

no

pooja195 (pooja195):

my bad 3+1=4 -2+4=2 4/2=2 :/

OpenStudy (anonymous):

So we are going to use the slope formula first so \[\large \frac{-1+(-3)}{4-(-2)}\]